# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v2

@@ IV-D Visualization at Distribution Level
Figure 8: t-SNE Visualization of Cross-modal Feature Distribution Evolution. (a) RGB encoder features show spectral continuity but blurred class boundaries. (b) SAR encoder features capture geometric details but are affected by speckle noise. (c) PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separation.
As shown in Fig. 8, PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability. Quantitatively, mean intra-class standard variance is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland, compared to the best unimodal baselines. Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% (66.56 $to$ 116.95) and Farmland/Forest by 13.29% (107.64 $to$ 121.89), indicating that PAD captures cross-modal discriminative patterns effectively.
TABLE IX: Statistical Comparison Between Models w/ and w/o ASF.
@@
Paired t-test p-value 0.0016
@@
IV-E Noise Sensitivity of ASF
To assess whether ASF amplifies high-frequency speckle noise in SAR images, we conducted a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle. For each image, we computed pixel-wise ENL maps and then measured the Pearson and Spearman correlations between ENL and classification accuracy. As shown in Table IX, the model with ASF exhibits nearly identical correlation coefficients to the baseline, differences in mean correlation are below 0.001, and Cohen’s $d=-0.038$ indicates a negligible effect size. These results confirm that ASF does not increase noise sensitivity and remains robust in low-SNR regions.
Overall, these results confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
@@ V Conclusion
In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification. Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, while amplitude spectra capture modality-specific details, providing the first physical prior for multi-source remote sensing fusion. PAD disentangles and aligns phase features via the Phase Spectrum Correction (PSC) module, and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, which integrates SAR’s high-frequency details with RGB’s low-frequency context. Extensive experiments on two challenging datasets show that PAD outperforms existing methods in both accuracy and robustness.

### v3

@@ IV-D1 Visualization at the Distribution Level
Figure 8: t-SNE Visualization of Cross-Modal Feature Distribution Evolution. (a) RGB-encoder features show spectral continuity but exhibit blurred class boundaries; (b) SAR-encoder features capture geometric details but are affected by speckle noise; (c) PAD-fused features demonstrate enhanced intra-class compactness and improved inter-class separability.
As shown in Fig. 8, PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability. Quantitatively, mean intra-class standard deviation is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland compared with the best unimodal baselines. Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% ($66.56\rightarrow 116.95$) and Farmland/Forest by 13.29% ($107.64\rightarrow 121.89$), indicating that PAD captures cross-modal discriminative patterns effectively.
TABLE IX: Performance Comparison (%) of Models w/ and w/o PSC under Missing RGB Regions and Cloud Obscuration on WHU-OPT-SAR.
Scenario Method Per-class IoU OA Kappa mF1 mIoU
Farmland City Village Water Forest Road Others
Missing RGB w/o PSC 74.65 75.91 58.03 80.18 56.84 57.81 59.49 82.28 70.94 79.23 66.13
w/ PSC 74.65 76.25 58.05 80.25 57.09 58.54 59.82 82.35 71.12 79.42 66.38
Cloud Obscuration w/o PSC 75.33 75.47 57.42 76.31 84.30 53.40 44.94 86.90 63.06 79.23 66.74
w/ PSC 75.37 75.73 57.56 76.47 84.32 53.68 45.30 86.95 63.52 79.37 66.92
IV-D2 Robustness of PSC under Challenging Scenarios
To further evaluate the robustness of PSC beyond normal distributions, we conducted targeted ablation experiments under extreme long-tail conditions, including missing RGB regions and cloud obscuration. Results are summarized in Table IX.
Missing Regions
When large portions of the RGB modality are absent, models without PSC accumulate residual phase errors, particularly in categories with fine spatial structures (e.g., forest, road). With PSC, alignment remains stable and yields consistent improvements across macro-level metrics (e.g., +0.19 mF1, +0.25 mIoU).
Cloud Obscuration
Under cloud occlusion, traditional co-registration often provides weak or misleading guidance. PSC mitigates this issue by preserving phase consistency, resulting in systematic improvements in category-wise IoU (e.g., city, village) as well as overall agreement scores.
These results confirm that PSC not only aligns with physical and statistical expectations but also maintains robust performance under adverse long-tail scenarios, where purely physical models typically degrade.
TABLE X: Statistical Comparison between Models w/ and w/o ASF.
@@ IV-D2 Robustness of PSC under Challenging Scenarios
Paired $t$-test pp-value 0.0016
@@
IV-D3 Noise Sensitivity of ASF
To assess whether ASF amplifies high-frequency speckle noise in SAR images, we perform a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle. For each image, we compute pixel-wise ENL maps and then measure Pearson and Spearman correlations between ENL and classification accuracy. As shown in Table X, the model with ASF exhibits nearly identical correlation coefficients to the baseline; differences in mean correlation are below 0.001, and Cohen’s d=0.038d=-0.038 indicates a negligible effect size. These results confirm that ASF does not increase noise sensitivity and remains robust in low-SNR regions.
Overall, these findings confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
@@ V Conclusion
In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land-cover classification. Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, whereas amplitude spectra capture modality-complementary details, providing the first physically grounded prior for multi-source remote-sensing fusion. PAD decouples and aligns phase features via the Phase Spectrum Correction (PSC) module and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, effectively leveraging SAR’s high-frequency details together with RGB’s low-frequency context. Extensive experiments on two challenging datasets show that PAD outperforms existing methods in both accuracy and robustness.